### 08 Apr 2024 v2

@@ 2.1 LLM with Chinese Language Ability
In the field of LLMs, the advancement of technologies has catalyzed the development of an array of open-source models exhibiting remarkable linguistic capabilities. Notably, models such as LLaMA Touvron et al. (2023a; b), Phi Li et al. (2023b); Gunasekar et al. (2023), Mistral Jiang et al. (2023), and Gemma Team et al. (2024) have emerged as frontrunners, underscoring the technological strides made in this arena. Amidst a globalized context, there’s a rising demand for models proficient in bilingual or multilingual functionalities, particularly those accommodating the vast spectrum of Chinese language applications. This demand stems from the desire for localized solutions and the necessity to bridge linguistic divides worldwide. To address this need, several strategies have been employed to enhance the multilingual capabilities of LLMs, with a significant emphasis on incorporating a higher proportion of Chinese tokens during the pretraining phase or employing techniques such as supervised fine-tuning (SFT) to activate Chinese language functionalities Zeng et al. (2023); Bai et al. (2023); Yang et al. (2023); Team (2023); Young et al. (2024); Bi et al. (2024). An early example in this endeavor is ChatGLM Zeng et al. (2023), which pioneered the use of an equal distribution of Chinese and English tokens during its pretraining phase, culminating in a proficient bilingual model. Following this, models like Qwen Bai et al. (2023) have expanded the linguistic horizon by integrating multilingual data in the pretraining process, thereby achieving broader language support. Furthermore, models such as Yi Young et al. (2024) and DeepSeek Bi et al. (2024) have demonstrated the efficacy of meticulous SFT applications in unlocking multilingual capabilities, with a notable prowess in Chinese language reasoning. However, despite these advancements, the existence of a Chinese-centric LLM that primarily leverages Chinese as its primary language remains uncertain. This gap highlights a critical area of interest for developing localized, open-source Chinese models, underscoring the significance of tailored approaches in the evolution of language technologies.
@@ 2.3 Emergence of Multilingual Capacity
The prevailing paradigm in developing LLMs has largely favored English-centric pretraining methodologies. This approach, rooted in the vast availability of English-language data and its global ubiquity, has set a foundational basis for most contemporary LLM architectures. Subsequently, strategies such as continuing pretraining, supervised fine-tuning, and instruction fine-tuning (IFT) have been employed to extend these models’ linguistic reach, enabling the activation of multilingual capacities Zeng et al. (2023); Bai et al. (2023); Yang et al. (2023); Team (2023); Young et al. (2024); Bi et al. (2024). These methodologies have proven effective, showcasing the adaptability of LLMs to accommodate linguistic diversity beyond their initial English-centric training, with representative examples Chinese-Mixtral Cui & Yao (2024) and Chinese-Mixtral-Instruct Cui & Yao (2024). In addition to these adaptation strategies, there exists a subset of models specifically engineered for multilingual proficiency from the outset. Models like BLOOM Le Scao et al. (2022) and Aya Üstün et al. (2024) exemplify this approach, incorporating a multitude of languages throughout both their pretraining and fine-tuning phases. Despite these efforts to integrate linguistic diversity, English invariably remains the dominant language within these models Zhao et al. (2024). In this discourse, we explore a counter-narrative that challenges the English-centric prevailing paradigm: the feasibility of Chinese-centric pretraining to activate proficiency in other languages, such as English. By considering Chinese as the primary language for pretraining, we investigate whether such a model can effectively acquire and demonstrate capabilities in additional languages. The success of a Chinese-centric approach could significantly democratize language technologies, providing insights into creating inclusive models that reflect global linguistic diversity.
@@ 3.2 Model Architecture
Our model’s architecture is based on the transformer decoder Vaswani et al. (2017a). The key parameters that define our architecture are shown in Table 1, with the models being trained on a substantial context length of 4096 tokens. Beyond the foundational elements, our approach integrates several improvements compared to the original transformer.
Multi-Head Attention Mechanism. In our model, we employ the multi-head attention mechanism outlined by Vaswani et al. (2017b). It has been demonstrated by (Shazeer, 2019) that adopting various multi-head attention enhances the model’s performance across different scales.

### 09 Apr 2024 v3

@@ 2.1 LLM with Chinese Language Ability
In the field of LLMs, the advancement of technologies has catalyzed the development of an array of open-source models exhibiting remarkable linguistic capabilities. Notably, models such as LLaMA (Touvron et al., 2023a; b), Phi (Li et al., 2023b; Gunasekar et al., 2023), Mistral (Jiang et al., 2023), and Gemma (Team et al., 2024) have emerged as frontrunners, underscoring the technological strides made in this arena. Amidst a globalized context, there’s a rising demand for models proficient in bilingual or multilingual functionalities, particularly those accommodating the vast spectrum of Chinese language applications. This demand stems from the desire for localized solutions and the necessity to bridge linguistic divides worldwide. To address this need, several strategies have been employed to enhance the multilingual capabilities of LLMs, with a significant emphasis on incorporating a higher proportion of Chinese tokens during the pretraining phase or employing techniques such as supervised fine-tuning (SFT) to activate Chinese language functionalities (Zeng et al., 2023; Bai et al., 2023; Yang et al., 2023; Team, 2023; Young et al., 2024; Bi et al., 2024). An early example in this endeavor is ChatGLM (Zeng et al., 2023), which pioneered the use of an equal distribution of Chinese and English tokens during its pretraining phase, culminating in a proficient bilingual model. Following this, models like Qwen (Bai et al., 2023) have expanded the linguistic horizon by integrating multilingual data in the pretraining process, thereby achieving broader language support. Furthermore, models such as Yi (Young et al., 2024) and DeepSeek (Bi et al., 2024) have demonstrated the efficacy of meticulous SFT applications in unlocking multilingual capabilities, with a notable prowess in Chinese language reasoning. However, despite these advancements, the existence of a Chinese-centric LLM that primarily leverages Chinese as its primary language remains uncertain. This gap highlights a critical area of interest for developing localized, open-source Chinese models, underscoring the significance of tailored approaches in the evolution of language technologies.
@@ 2.3 Emergence of Multilingual Capacity
The prevailing paradigm in developing LLMs has largely favored English-centric pretraining methodologies. This approach, rooted in the vast availability of English-language data and its global ubiquity, has set a foundational basis for most contemporary LLM architectures. Subsequently, strategies such as continuing pretraining, supervised fine-tuning, and instruction fine-tuning (IFT) have been employed to extend these models’ linguistic reach, enabling the activation of multilingual capacities (Zeng et al., 2023; Bai et al., 2023; Yang et al., 2023; Team, 2023; Young et al., 2024; Bi et al., 2024). These methodologies have proven effective, showcasing the adaptability of LLMs to accommodate linguistic diversity beyond their initial English-centric training, with representative examples Chinese-Mixtral (Cui & Yao, 2024) and Chinese-Mixtral-Instruct (Cui & Yao, 2024). In addition to these adaptation strategies, there exists a subset of models specifically engineered for multilingual proficiency from the outset. Models like BLOOM (Le Scao et al., 2022) and Aya (Üstün et al., 2024) exemplify this approach, incorporating a multitude of languages throughout both their pretraining and fine-tuning phases. Despite these efforts to integrate linguistic diversity, English invariably remains the dominant language within these models (Zhao et al., 2024). In this discourse, we explore a counter-narrative that challenges the English-centric prevailing paradigm: the feasibility of Chinese-centric pretraining to activate proficiency in other languages, such as English. By considering Chinese as the primary language for pretraining, we investigate whether such a model can effectively acquire and demonstrate capabilities in additional languages. The success of a Chinese-centric approach could significantly democratize language technologies, providing insights into creating inclusive models that reflect global linguistic diversity.
@@ 3.2 Model Architecture
Our model’s architecture is based on the transformer decoder (Vaswani et al., 2017a). The key parameters that define our architecture are shown in Table 1, with the models being trained on a substantial context length of 4096 tokens. Beyond the foundational elements, our approach integrates several improvements compared to the original transformer.
Multi-Head Attention Mechanism. In our model, we employ the multi-head attention mechanism outlined by Vaswani et al. (2017b). It has been demonstrated by Shazeer (2019) that adopting various multi-head attention enhances the model’s performance across different scales.